# An End-to-End Depth-Based Pipeline for Selfie Image Rectification

### v1

@@ 5.3 Results
TABLE II: Quantitative comparison on CMDP dataset [31]. LMKE is the landmark distance error and refers to the Euclidean distance between the corresponding facial landmarks. LMKE is normalized by the image size which is $512\times 512$
@@
Input Ours Fried [6] Shih [4] Disco [7] Reference
Figure 6: Qualitative comparison on the CMDP dataset [31].
We also conduct a comparison on the Caltech Multi-Distance Portraits (CMDP) dataset [31]. This dataset contains portrait images of several subjects captured from various distances. For each identity, the camera-to-subject distance is varied as well. Since the ground truth relative poses between the cameras are not known and the images for the same persons were not synchronized, it is hard to use direct intensity-based metrics such as PSNR or SSIM for measuring the quality. Instead, we estimate the Euclidean distance errors between the aligned facial landmarks of the output image and the pseudo-target image. We use the images captured at distances of 60cm and 480cm as the input and reference, respectively. Using the facial landmarks, we estimate a similarity transform to align the predicted and reference images and compute the perceptual similarity LPIPS on the masked foreground. Table II, shows a quantitative comparison on the CMDP dataset. Additionally, Figure 6 shows a qualitative comparison on the same dataset. Our visual results look more consistent with the reference image than the correction of Fried [6] and Shih [4], while being comparable with the results of the time-consuming Disco method [7].
@@
Furthermore, Figure 8 shows some correction result comparison on in-the-wild images collected by Disco [7]. Again, our method’s rectification looks more faithful than Fried [6], and is quite comparable with Disco. It is important to note that we don’t know the focal length of the cameras (or the field of view) in these cases, but it is likely to be quite different from the case used in generating the training data. Despite this, the method works well and clearly generalizes not only to real images but also to other configurations.
@@
Finally, we compare the full-frame rectification quality with the Disco method [7]. Figure 9, shows a qualitative comparison of full-frame results between our pipeline and Disco [7]. Our method produces similar final results as Disco without separate processing of face crops.
@@
Figure 9: Full-frame image rectification comparison. The last image is a screenshot from the result website provided by Disco method [7]
5.4 Ablation studies
@@
Figure 11: Qualitative comparison on the synthetic dataset when disabling the horizontal translation module (HzT). The predicted output and GT target are shown in both cases when $tx=0$ and the $t_{x}$ predicted by our model. The PSNR is reported on the right side of the image while the predicted $t_{x}$ is shown on the left.
@@
Input $t_{x}=0$ $t_{x}\neq 0$
Figure 12: Qualitative comparison on the real images illustrating the effect of ablating the horizontal translation module. The orientation when $t_{x}\neq 0$ gives the sense that there are no missing face parts such as ears. Additionally, the head orientation conforms with its corresponding input.
Next, we demonstrate the benefit of incorporating the auxiliary translation module to enhance the synthesized novel views. Figure 11 shows a comparison on the synthetic dataset when we ablate the translation module from our pipeline. The module can effectively predict a translation that avoids the hard missing regions in the face boosting the quality of the synthesized novel views in those cases. In Figure 12, we show the results of ablating the horizontal translation module on real data. Translating the camera horizontally gives the sense that there are no missing parts of the face such as the ear. In addition, the head orientation after translating the camera conforms with the orientation in its corresponding input. To demonstrate the generalizability of our modules trained on a synthetic dataset, we show images captured by different mobile phone cameras that have different focal lengths. The images of Figure 12 are captured by Samsung S23 Ultra, Huwaei P40Pro, and iPhone 12, respectively. Their 35mm-equivalent focal length are: 25, 26, and 23, and their field of view angles are: 81.6°, 79.5°, and 86.5°. Our pipeline can deal with subjects wearing eyeglasses as shown in the third row of Figure 12.
@@
6 Limitations and Future Work

### v2

@@ V-C Results
TABLE II: Quantitative comparison on the CMDP dataset [33]. LMKE denotes the landmark error, normalized by the image size ($512\times 512$ pixels). Methods marked with (*) are evaluated on a smaller subset of the CMDP dataset used in [14].
@@
DFN* [14] 0.00613 0.2628
Ours* 0.00499 0.2205
Input Ours Fried [6] Shih [4] DP [32] Disco [7] Reference
Figure 6: Qualitative comparison on the CMDP dataset [33]. The DP method [32] produces an image and a mask. The mask can be used to fill the background holes with an external inpainting method. Here, we show the direct output obtained by running their code.
We also conduct a comparison on the Caltech Multi-Distance Portraits (CMDP) dataset [33]. This dataset contains portrait images of several subjects captured from various distances. For each identity, the camera-to-subject distance is varied as well. Since the ground truth relative poses between the cameras are not known and the images for the same persons were not synchronized, it is hard to use direct intensity-based metrics such as PSNR or SSIM for measuring the quality. Instead, we estimate the Euclidean distance errors between the aligned facial landmarks of the output image and the pseudo-target image. We use the images captured at distances of 60cm and 480cm as the input and reference, respectively. Using the facial landmarks, we estimate a similarity transform to align the predicted and reference images and compute the perceptual similarity LPIPS on the masked foreground. Table II, shows a quantitative comparison on the CMDP dataset. Additionally, Figure 6 shows a qualitative comparison on the same dataset. Our visual results look more consistent with the reference image than the correction of Fried [6], Shih [4] and DP [32], while being comparable with the results of the time-consuming Disco method [7]. Note that Shih [4] and DP [32] are not able to rectify perspective distortion in close-range portraits.
@@
Furthermore, Figure 8 shows some correction result comparisons on in-the-wild images collected by Disco [7]. Again, our method’s rectification looks more faithful than Fried [6], and is quite comparable with Disco. It is important to note that we don’t know the focal length of the cameras (or the field of view) in these cases, but it is likely to be quite different from the case used in generating the training data. Despite this, the method works well and clearly generalizes not only to real images but also to other configurations.
@@
Finally, we compare the full-frame rectification quality with the Disco method [7]. Figure 9 shows a qualitative comparison of full-frame results between our pipeline and Disco [7]. Our method produces similar final results as Disco without separate processing of face crops.
@@
Figure 9: Full-frame image rectification comparison. The last image is a screenshot from the result website provided by the Disco method [7].
V-D Ablation studies
@@
Figure 11: Qualitative comparison on the synthetic dataset when disabling the horizontal translation module (HzT). The predicted output and GT target are shown in both cases when $t_{x}=0$ and the $t_{x}$ predicted by our model. The PSNR is reported on the right side of the image while the predicted $t_{x}$ is shown on the left.
@@
Input $t_{x}=0$ $t_{x}\neq 0$ Inpainted region
Figure 12: Qualitative comparison on the real images illustrating the effect of ablating the horizontal translation module. The orientation when $t_{x}\neq 0$ gives the sense that there are no missing face parts such as ears. Additionally, the head orientation conforms with its corresponding input. The last two columns visualize the inpainted regions, highlighted in red, for each case. The percentage of the inpainted area is indicated on the left, while the reduction in inpainted area (when using horizontal translation) is shown in the top-right corner of each image.
Next, we demonstrate the benefit of incorporating the auxiliary translation module to enhance the synthesized novel views. Figure 11 shows a comparison on the synthetic dataset when we ablate the translation module from our pipeline. The module can effectively predict a translation that avoids the hard missing regions in the face boosting the quality of the synthesized novel views in those cases. In Figure 12, we show the results of ablating the horizontal translation module on real data. Translating the camera horizontally gives the sense that there are no missing parts of the face such as the ear. In addition, the head orientation after translating the camera conforms with the orientation in its corresponding input. We also visualize the inpainted area for each case in the last two columns of Figure 12. It is shown that the inpainted region in the case of $t_{x}\neq 0$ is smaller. To demonstrate the generalizability of our modules trained on a synthetic dataset, we show images captured by different mobile phone cameras that have different focal lengths. The images of Figure 12 are captured by Samsung S23 Ultra, Huawei P40Pro, and iPhone 12, respectively. Their 35mm-equivalent focal length are: 25, 26, and 23, and their field of view angles are: 81.6°, 79.5°, and 86.5°. Our pipeline can deal with subjects wearing eyeglasses as shown in the third row of Figure 12.
@@
VI Limitations and Future Work